Four resistors when connected in series give an equivalent


resistance if 75.0 Ω. Three of the resistors have values 18.0 Ω, 6.0 Ω, and


24.0 Ω, respectively. What is the resistance of the fourth resistor?

1
Expert's answer
2022-03-23T13:53:11-0400

Answer

the resistance of the fourth resistor is given by

"R_4=R-R_1-R_2-R_3\\=75-(18+6+24) \\=27\Omega"